xx HTTPDExtra

HTTPDExtra

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11viewportxrefreshxseconds 10Scope: privatex01  INT

11Description: Viewport web page refresh time seconds

11 Range 10 x01 Default: -1
11-1:  10-1 for no refresh, 0 for immediate refresh

11

7 Interfaces

General Implements: httpxutils

Uses header: httpxAuth.h httpxContent.h

8 Schedule This section lists all the variables which are assigned storage by thorn CactusConnect/HTTPDExtra. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x httputilsxstartup x utils for httpd startup x

11x11After: 11httpxstartup
11 x 11Before: 11ioutilxstartup
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function